# Evaluating NeRFs for 3D Plant Geometry Reconstruction in Field Conditions

### v2

@@ 3.2. Scenario II - Multiple Plants Indoors
The estimated camera poses from colmap are visualized in Fig. 12. We observe marked differences in model behaviors compared to the single plant scenario, which are likely attributed to the added intricacy of multiple plants in a single scene. Results are illustrated in the Fig. 13.
Figure 12. Camera Pose Estimations for Scenario-II
(a) PSNR
(b) SSIM
(c) LPIPS
(d) Precision
(e) Recall
(f) F1 Score
Number of Iterations
  Instant-NGP    NeRFacto    TensoRF
Figure 13. Combined Comparison of 2D Image Quality (top) and 3D Geometry (bottom) Metrics for Scenario-II
@@
Precision: Instant-NGP shows a steady increase in precision with more iterations, starting from 0.57 at 100 iterations and peaking at 23.45 at 30000 iterations. However, NeRFacto starts at a higher precision of 7.77 at 100 iterations and reaches a higher peak of 64.47 at 30000 iterations, indicating a more accurate reconstruction of the corn plants. TensoRF, although it improves with more iterations, lags behind the others, starting at 1.51 and reaching 20.50 at 30000 iterations. It is further illustrated in Fig. 14.
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Figure 14. Precision of 3D reconstruction of NeRF techniques - Scenario II
Recall:
@@
For recall, a similar pattern is observed. Instant-NGP’s recall increases from 1.23 to 58.57, NeRFacto from 27.44 to 76.80, and TensoRF from 12.62 to 55.34 across iterations. NeRFacto consistently maintains a higher recall, suggesting its superior ability to encompass points in the ground truth.. It is further illustrated in Fig. 15.
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Figure 15. Recall of 3D reconstruction of NeRF techniques - Scenario II
F1 Score:
@@
The F1 Score, balancing precision and recall, also follows this trend. Instant-NGP’s F1 score peaks at 33.49, NeRFacto at 70.10, and TensoRF at 29.91, all at 30000 iterations. NeRFacto demonstrates the best balance between precision and recall.
PSNR:
In terms of PSNR, which evaluates the quality of rendered images, all models show improvement with more iterations. Instant-NGP goes from 13.70 to 19.08, NeRFacto from 14.93 to 18.93, and TensoRF from 13.38 to 15.54. Instant-NGP achieves the highest PSNR, suggesting better image quality.
SSIM:
For SSIM, higher values indicate better image structure similarity. Instant-NGP progresses from 0.36 to 0.64, NeRFacto from 0.35 to 0.64, and TensoRF from 0.35 to 0.42. Both Instant-NGP and NeRFacto perform similarly and better than TensoRF in this aspect.
LPIPS:
Lower LPIPS values signify higher perceptual similarity to the ground truth. Instant-NGP decreases from 0.89 to 0.31, NeRFacto from 0.77 to 0.25, and TensoRF from 0.83 to 0.56, with NeRFacto showing the best perceptual image quality.
Computation Time:
@@
The time taken for iterations is crucial for efficiency. Instant-NGP and NeRFacto have comparable times, but TensoRF takes significantly longer at higher iterations, indicating less time efficiency.
Overall Performance and Suitability:
@@
NeRFacto emerges as the most balanced and efficient model, exhibiting high precision, recall, and F1 scores, along with favorable PSNR, SSIM, and LPIPS values. Its efficiency in time taken is also comparable to Instant-NGP. Instant-NGP, while showing improvements, doesn’t quite match NeRFacto’s balance of precision and recall. TensoRF, despite its merits, falls behind in several key metrics, particularly in precision, recall, SSIM, and LPIPS. Detailed results are given in Table 4, after complete training. For more granular look, consult the supplementary Table 7. The Precision-Recall Curves based on varying distance threshold after maximum training of 30,000 iterations is given in Fig. 16.
Table 4. Performance metrics of NeRFs reconstruction techniques - Scenario II
Model Name Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time(s) $\downarrow$
Instant-NGP 23.45 58.57 33.49 19.08 0.64 0.31 1886
NeRFacto 64.47 76.8 70.1 18.93 0.64 0.25 1226
TensoRF 20.5 55.34 29.91 15.54 0.42 0.56 2607
# of points (%)
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Meters
  Precision    Recall
Figure 16. Precision-Recall Curves (Scenario-II) based on varying distance threshold after 30,000 iterations
Insight 1: Improved Performance of TensoRF in Scenario 2:
@@
In the second scenario, TensoRF demonstrated a notable improvement compared to its performance in the first scenario. Specifically, its F1 score, a critical metric for 3D modeling accuracy, significantly increased from 15.69 in the first scenario to 29.91 after 30,000 iterations in the second scenario. This substantial improvement highlights TensoRF’s potential in more complex or demanding 3D modeling tasks, especially when allowed to complete its training process.
Insight 2: 2D Metrics Versus 3D F1 Score for Instant-NGP and NeRFacto:
@@
While Instant-NGP and NeRFacto show comparable results in 2D image quality metrics such as PSNR and SSIM, a distinct difference is observed in their 3D modeling capabilities, as reflected in their F1 scores, as observed in last scenario. This suggests that NeRFacto might be a more reliable choice for applications requiring high accuracy in 3D reconstructions.
3.3. Scenario III - Multiple Plants Outdoors
The estimated camera poses from colmap are visualized in Fig. 17
Figure 17. Camera Pose Estimations for Scenario-III
The results from our evaluation of three state-of-the-art NeRF techniques for 3D plant structure reconstruction are illustrated in the Fig. 18. Please note that the value of threshold distance $d$ for 3d metrics was considered 0.01, as compared to 0.005 in last two scenarios.
(a) PSNR
(b) SSIM
(c) LPIPS
(d) Precision
(e) Recall
(f) F1 Score
Number of Iterations
  Instant-NGP    NeRFacto    TensoRF
Figure 18. Combined Comparison of 2D Image Quality (top) and 3D Geometry (bottom) Metrics for Scenario-III
Precision:
@@ 3.3. Scenario III - Multiple Plants Outdoors
Across all iterations, NeRFacto consistently demonstrates the highest precision, peaking at 68.29% at 60000 iterations, suggesting its superior ability to reconstruct points close to the ground truth. Instant-NGP shows a steady increase in precision with more iterations, yet never surpasses NeRFacto, reaching a maximum of 15.06%. TensoRF, while starting lower, reaches a comparable precision to Instant-NGP at higher iterations (40.95% at 60000 iterations). It is further illustrated in Fig. 19.
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Figure 19. Precision of 3D reconstruction of NeRF techniques - Scenario II
Recall:
@@
NeRFacto also leads in recall, achieving a high of 82.32% at 60000 iterations, indicating its effectiveness in encompassing points from the ground truth. Instant-NGP shows significant improvement with increased iterations, reaching a recall of 59.55%, but remains behind NeRFacto. TensoRF shows substantial growth in recall, ultimately achieving 75.62%, positioning it between Instant-NGP and NeRFacto in terms of completeness. It is further illustrated in Fig. 20.
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Figure 20. Recall of 3D reconstruction of NeRF techniques - Scenario II
F1 Score:
@@
Reflecting the balance between precision and recall, the F1 scores show NeRFacto as the superior model, with a peak score of 74.65% at 60000 iterations. Instant-NGP’s F1 score also improves with more iterations but tops out at 24.04%, significantly lower than NeRFacto. TensoRF’s F1 score, while lower than NeRFacto’s, surpasses Instant-NGP, reaching 53.13% at 60000 iterations.
PSNR:
In terms of image quality, as measured by PSNR, all models show incremental improvements with more iterations. Instant-NGP and NeRFacto display similar trends, with NeRFacto slightly leading, peaking at 16.70 at 60000 iterations. TensoRF shows a comparable maximum PSNR of 17.32, suggesting its slight edge in rendering higher-quality images.
SSIM:
For the SSIM metric, all three models show improvements with increased iterations. NeRFacto maintains a slight advantage over the others, peaking at 0.32 at 60000 iterations, indicating its better performance in maintaining structural integrity in the rendered images. Instant-NGP and TensoRF show similar SSIM scores, with TensoRF slightly leading at higher iterations.
LPIPS:
The LPIPS scores, which assess perceptual similarity, decrease for all models with more iterations, indicating improved performance. NeRFacto and TensoRF show similar trends, with NeRFacto having a slight edge, achieving a score of 0.34 at 60000 iterations compared to TensoRF’s 0.55. Instant-NGP’s performance is consistently lower in this metric.
Computation Time:
@@
In terms of efficiency, Instant-NGP and NeRFacto are the fastest, followed by TensoRF.
Overall Performance and Suitability:
@@
NeRFacto again emerges as the most balanced and robust model, excelling in precision, recall, F1 score, and LPIPS. Detailed results are given in Table 5, after complete training. For more granular look, consult the supplementary Table 8. The Precision-Recall Curves based on varying distance threshold after maximum training of 60,000 iterations is given in Fig. 21.
Table 5. Performance metrics of NeRFs reconstruction techniques - Scenario III
Model Name Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time(s) $\downarrow$
Instant-NGP 15.06 59.55 24.04 18.54 0.47 0.4 1466
NeRFacto 68.29 82.32 74.65 16.7 0.32 0.34 1499
TensoRF 40.95 75.62 53.13 17.32 0.39 0.55 1965
# of points (%)
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Meters
  Precision    Recall
Figure 21. Precision-Recall Curves (Scenario-III) based on varying distance threshold after 60,000 iterations
Insight 1: Enhanced Performance of TensoRF in Outdoor Settings:
@@
TensoRF demonstrates significant improvement in its performance in the third scenario compared to the first. Specifically, its F1 score has seen a notable increase; from 15.69 in the first scenario to 29.91 in the second, and reaching 53.13 after 30,000 iterations in the current outdoor scenario. This upward trajectory in F1 scores, which is a balanced measure of precision and recall, indicates TensoRF’s enhanced capability in outdoor environments, potentially outperforming Instant-NGP in these settings. This suggests that TensoRF might be a more suitable choice for outdoor 3D modeling tasks where both precision and completeness are crucial. This property may have contributed in the selection of TensoRF as a building block for using multiple local radiance fields, during in-the-wild reconstruction, in the recent paper [Meuleman et al., 2023].
Insight 2: LPIPS as a Strong Indicator of 3D Model Quality:
@@
The LPIPS metric appears to be a more representative measure of the quality of the resulting 3D models. In the analysis, we observe that models with lower LPIPS scores consistently show better performance across other metrics. This trend indicates the relevance of LPIPS in assessing the perceptual quality of 3D models. The further investigation into how LPIPS correlates with other metrics could provide deeper insights into model performance, especially in the context of realistic and perceptually accurate 3D reconstructions.
4. Discussion
@@
The difference in the output quality between Instant-NGP and Nerfacto, especially concerning the density and crispness of the rendered scenes, could indeed be related to the sampling strategies used by each algorithm.
Instant-NGP Sampling Strategy:
@@
Instant-NGP uses an improved training and rendering algorithm that involves a ray marching scheme with an occupancy grid. This means that when the algorithm shoots rays into the scene to sample colors and densities, it uses an occupancy grid to skip over empty space, as well as areas behind high-density regions to improve efficiency.
  • The occupancy grid used in Instant-NGP is a multiscale grid that coarsely marks empty and non-empty space and is used to determine where to skip samples to speed up processing.
  • This approach is quite effective in terms of speed, leading to significant improvements over naive sampling methods.
@@
  • However, if the occupancy grid isn’t fine-grained enough or if the method for updating this grid isn’t capturing the scene’s density variations accurately, it could lead to a "muddy" or overly dense rendering because it might not be sampling the necessary areas with enough precision.
Nerfacto Sampling Strategy:
@@
Nerfacto, on the other hand, uses a combination of different sampling techniques:

### v3

@@ 3.2   Scenario II - Multiple Plants Indoors
We observe marked differences in model behaviors compared to the single plant scenario, likely attributed to the added intricacy of multiple plants in a single scene. Detailed evolution of each metric over training iterations is given in the Supplement.
@@
Precision: As shown in Figure 6, Instant-NGP exhibits a steady increase in precision with more iterations, peaking at a high value. However, NeRFacto starts at a higher precision and reaches an even higher peak, indicating a more accurate reconstruction of the corn plants. TensoRF, although improving with more iterations, lags behind the others in terms of precision.
@@
Recall: A similar pattern is observed for recall, with NeRFacto consistently maintaining a higher recall compared to the other methods, suggesting its ability to better encompass points in the ground truth. Both Instant-NGP and TensoRF exhibit increasing recall with more iterations, but at lower levels than NeRFacto.
@@
F1 Score: The F1 Score, balancing precision and recall, follows a similar trend. NeRFacto demonstrates the best balance between precision and recall, with its F1 score peaking at 70.10, while Instant-NGP and TensoRF achieve lower peak F1 scores.
@@
Computation Time: The time taken for iterations is crucial for efficiency. Instant-NGP and NeRFacto have comparable times, but TensoRF takes significantly longer at higher iterations, indicating less time efficiency. See supplementary material for evolution of precision metric over the training iterations.
@@
Overall Performance and Suitability: NeRFacto emerges as the most balanced and efficient model, exhibiting high precision, recall, and F1 scores, along with favorable PSNR, SSIM, and LPIPS values. Its efficiency in time taken is also comparable to Instant-NGP. Instant-NGP, while showing improvements, doesn’t quite match NeRFacto’s balance of precision and recall. TensoRF, despite its merits, falls behind in several key metrics, particularly in precision, recall, SSIM, and LPIPS. The results after complete training are given in Table 2. For more granular look of each metric value over the training iteration for all the algorithms, consult the supplementary. The Precision-Recall curves based on varying distance thresholds after maximum training of 30,000 iterations are given in Figure 7.
@@
Insight 1: Improved Performance of TensoRF in Scenario II: In the second scenario, TensoRF demonstrated an improvement compared to its performance in the first scenario. Specifically, its F1 score, a critical metric for 3D modeling accuracy, increased from 15.69 in the first scenario to 29.91 after 30,000 iterations in the second scenario. This improvement highlights TensoRF’s potential in more complex or demanding 3D modeling tasks, especially when allowed to complete its training process.
@@
Insight 2: 2D Metrics Versus 3D F1 Score for Instant-NGP and NeRFacto: While Instant-NGP and NeRFacto show comparable results in 2D image quality metrics such as PSNR and SSIM, a distinct difference is observed in their 3D modeling capabilities, as reflected in their F1 scores, as observed in last scenario. This suggests that NeRFacto might be a more reliable choice for applications requiring high accuracy in 3D reconstructions.
3.3   Scenario III - Multiple Plants Outdoors
Scenario III is the most complex, with multiple overlapping plants captured in field conditions. The models were also trained until 60,000 iterations, while the previous two scenarios were trained only for 30,000 iterations. Detailed evolution of each metric over training iterations is given in the Supplement.
@@ 3.3   Scenario III - Multiple Plants Outdoors
Precision: As observed in Figure 6, NeRFacto consistently demonstrates the highest precision across all iterations, peaking at 68.29%, suggesting its ability to reconstruct points close to the ground truth. Instant-NGP shows a steady increase in precision with more iterations, while TensoRF, although starting lower, reaches a comparable precision to Instant-NGP at higher iterations.
@@
Recall: NeRFacto leads in recall, achieving a high of 82.32%, indicating its effectiveness in encompassing points from the ground truth. Instant-NGP shows significant improvement in recall with increased iterations, but remains behind NeRFacto. TensoRF’s recall growth positions it between Instant-NGP and NeRFacto in terms of completeness.
@@
F1 Score: Reflecting the balance between precision and recall, NeRFacto emerges as the superior model, with its F1 score peaking at 74.65%. Instant-NGP’s F1 score improves with more iterations but remains significantly lower, while TensoRF’s F1 score surpasses Instant-NGP, reaching 53.13%.
@@
Computation Time: In terms of efficiency, Instant-NGP and NeRFacto are the fastest, followed by TensoRF.
@@
Overall Performance and Suitability: NeRFacto again emerges as the most balanced and robust model, excelling in precision, recall, F1 score, and LPIPS. Detailed results are given in Table 2, after complete training. For more granular look of each metric value over the training iteration for all the algorithms, consult the supplementary. The Precision-Recall curves based on varying distance threshold after maximum training of 60,000 iterations is given in Figure 7. The GPU memory usage of this scenario comes out to be approximately a constant 3GB (for the total memory of GPU being 80GB).
@@
Insight 1: Enhanced Performance of TensoRF in Outdoor Settings: TensoRF demonstrates significant improvement in its performance in the third scenario compared to the first. Specifically, its F1 score has seen a good increase; from 15.69 in the first scenario to 29.91 in the second, and reaching 53.13 after 30,000 iterations in the current outdoor scenario. This upward trajectory in F1 scores, which is a balanced measure of precision and recall, indicates TensoRF’s enhanced capability in outdoor environments, potentially outperforming Instant-NGP in these settings. This suggests that TensoRF might be a more suitable choice for outdoor 3D modeling tasks where both precision and completeness are crucial. This property may have contributed in the selection of TensoRF as a building block for using multiple local radiance fields, during in-the-wild reconstruction [54].
@@
Insight 2: LPIPS as a Strong Indicator of 3D Model Quality: The LPIPS metric appears to be a more representative measure of the quality of the resulting 3D models. In the analysis, we observe that models with lower LPIPS scores consistently show better performance across other metrics. This trend indicates the relevance of LPIPS in assessing the perceptual quality of 3D models. The further investigation into how LPIPS correlates with other metrics could provide deeper insights into model performance, especially in the context of realistic and perceptually accurate 3D reconstructions.
3.4   Early Stopping Algorithm
The implementation of early stopping based on the LP-IPS metric yielded substantial savings in computational time across all scenarios, with a minor sacrifice in the fidelity of 3D reconstructions, as measured by the F1 score. Time savings were notable across the three tested methodologies—Instant-NGP, TensorRF, and NeRFacto—with each showing a marked decrease in training time without a commensurate loss in F1 score accuracy. For a deeper look of LPIPS, F1 Score and the recommended stopping point for each case, consult the supplementary material.
On average, the early stopping strategy resulted in a 61.1% reduction in training time, suggesting a significant efficiency gain in the process of 3D plant reconstruction using neural radiance fields. Concurrently, the average F1 score loss was contained to 7.4%, indicating that the early plateau detection has a moderate impact on the quality of the 3D point cloud reconstructions. Specifically, Instant-NGP presented a more pronounced variation in F1 score loss, which was notably higher in Scene-III, thereby affecting its average loss more than TensorRF and NeRFacto. TensorRF and NeRFacto showed a remarkable consistency in time savings, which was mirrored in their comparable F1 score losses, highlighting the robustness of these methods in early stopping scenarios.
These findings articulate a compelling case for the utilization of early stopping in NeRF-based 3D reconstruction tasks, emphasizing the need to balance between computational resources and reconstruction precision. Such a balance is pivotal in scenarios where time efficiency is paramount yet a minimal compromise on reconstruction accuracy is permissible.
3.5   Scenario IV - Validation Examples In Field Conditions
Scene-1
Soybean
1000 Iterations
20000 Iterations*
60000 Iterations
Scene-2
Anthurium Hookeri
1000 Iterations
20000 Iterations*
60000 Iterations
Scene-3
Mixture of plants
1000 Iterations
20000 Iterations*
60000 Iterations
Scene-4
C. Floribundum
1000 Iterations
20000 Iterations*
60000 Iterations
Scene-5
H. Paniculata
1000 Iterations
30000 Iterations*
60000 Iterations
Figure 8: Scenes for validating the early stopping algorithm and their 3D reconstructions: original scenes in the first column, iterative reconstructions in the right column, and optimal iterations in the third column (*).
The efficacy of the LPIPS-based early stopping algorithm was validated using a diverse dataset comprising images from five different types of plants captured in both indoor and outdoor settings, as illustrated in Figure 8. The validation process employed a threshold $\theta$ set to 0.005 and a consistency length $C$ of 6, with the granularity of interpolation fixed at 1000, spanning a total of 60000 training iterations. For practical application, checkpoints, inherently exponential in nature, necessitated linear interpolation to facilitate algorithm execution. Figure 8 shows the rendered point clouds at three stages: after 1000 iterations, at the recommended early stopping iteration, and upon completing the full 60000 iterations of training. Each row of the figure corresponds to one of the five validation scenes, providing a qualitative comparative analysis.
Notably, for all indoor scenes, the algorithm recommended halting training at 20000 iterations, whereas for outdoor scenes, the suggestion extended to 30000 iterations. This distinction underscores the algorithm’s sensitivity to environmental variables affecting perceptual similarity metrics. The rendered point clouds, particularly at the early stopping points, exhibit minimal visual discrepancies when compared to those obtained after the full training duration. By reducing computational demands without much loss in fidelity, this approach is a cost-effective strategy for enhancing modeling throughput in precision agriculture and botanical research. We substantiate the hypothesis that LPIPS can serve as a reliable surrogate for direct F1 score estimation in the context of NeRF training. The algorithm’s ability to accurately predict optimal stopping points—balancing computational efficiency with reconstruction accuracy—presents a compelling case for its adoption in scenarios where resource conservation is paramount, yet quality cannot be entirely sacrificed.
4   Discussion
In this section, we discuss the findings from our comparative analysis of NeRF models for 3D plant reconstruction. The results indicate that the Nerfacto model achieved the best performance, and we explore the theoretical basis for its superiority by examining the sampling strategies employed by the different models. Understanding these strategies provides insights into why Nerfacto outperformed the other models in terms of reconstruction quality. In our experiments, we found that the Nerfacto model produced the highest quality 3D reconstructions compared to Instant-NGP and other NeRF models. To understand the theoretical basis for Nerfacto’s superior performance, in this section we take a deeper look at the sampling strategies used by Nerfacto and Instant-NGP and how they influence the visual quality and level of detail in the rendered scenes.
@@ 4   Discussion
The divergent performance of the NeRF models necessitates a deeper examination of their underlying sampling strategies and their influence on the quality of 3D reconstruction. The difference in the output quality between Instant-NGP and Nerfacto, especially concerning the density and crispness of the rendered scenes, could indeed be related to the sampling strategies used by each algorithm.
@@
Instant-NGP Sampling Strategy: Instant-NGP uses an improved training and rendering algorithm that involves a ray marching scheme with an occupancy grid. This means that when the algorithm shoots rays into the scene to sample colors and densities, it uses an occupancy grid to skip over empty space, as well as areas behind high-density regions to improve efficiency.
@@
The occupancy grid used in Instant-NGP is a multiscale grid that coarsely marks empty and non-empty space and is used to determine where to skip samples to speed up processing. This approach is quite effective in terms of speed, leading to significant improvements over naive sampling methods. However, if the occupancy grid isn’t fine-grained enough or if the method for updating this grid isn’t capturing the scene’s density variations accurately, it could lead to a “muddy” or overly dense rendering because it might not be sampling the necessary areas with enough precision.
@@
NeRFacto Sampling Strategy: Nerfacto, on the other hand, uses a combination of different sampling techniques: